# FinPos: A Position-Aware Trading Agent System for Real Financial Markets

### v1

@@ 4.3. Multi-Timescale Reward Design
$d_{t}$ and $q_{t}$ denote the direction and quantity of the purchasing decision at the current time step, respectively.
This reward design serves two purposes: 1. Encourage trend alignment: The agent receives positive reinforcement when its predicted direction matches the actual market movement. In this way, the agent is guided away from overfitting to daily fluctuations and instead learns to associate price trajectories with textual information and broader market trends, enabling more robust causal inference in decision-making. 2. Penalty for passivity: In our experiments, we find that during periods of high volatility, agents tend to maintain a inactive state. To address this, we introduce a quadratic penalty on $|M_{t}|$. whenever the position remains unchanged. This discourages missed opportunities in volatile markets while also preventing excessive trading in stable periods.
5. Experiments
5.1. Experimental Setup
5.1.1. Datasets
@@
Our research utilizes actual financial data sourced from publicly authoritative providers and encompasses various forms of market information. (1) Stock prices, obtained from Yahoo Finance, include daily open-high-low-close-volume (OHLCV) data. (2) Company news, retrieved using the Finnhub stock API, includes each article’s related company name, publication date, headline, and summary text, processed for sentiment and semantic relevance. (3) Macroeconomic news, also retrieved via the Finnhub stock API, contains publication dates, headlines, and summary texts for each article. (4) 10-Q (quarterly reports) and 10-K (annual reports), accessed through the U.S. Securities and Exchange Commission (SEC) EDGAR API. These documents provide financial information that publicly listed companies are required to disclose, and standardized into a daily time series format.
5.1.2. Comparative Methods
@@
To assess the effectiveness of FinPos, we conduct comparative experiments with four LLM-based trading agents: FinGPT Yang et al. (2023), FINMEM Yu et al. (2024a), FinAgent Zhang et al. (2024b), and FINCON Yu et al. (2024b); three deep reinforcement learning (DRL) methods (A2C, PPO, DQN); two rule-based methods (MACD, RSI); and two market baselines (Random, Buy-and-Hold). All compared models are evaluated using identical data splits and executed under the same market conditions.
5.1.3. Evaluation Metrics
@@
We evaluate performance using cumulative return (CR%), sharpe ratio (SR), and maximum drawdown (MDD%). CR reflects overall profitability, while SR measures the risk-adjusted return by showing how much excess return it generates per unit of total risk. MDD quantifies the worst-case loss over the trading horizon, which in the Position-Aware Trading Task directly reflects exposure and vulnerability under held positions. Since CR has been introduced earlier (see Eq. (2)), here we provide the formulas for SR and MDD as follows:
$\text{Sharpe Ratio}=\frac{R_{p}-R_{f}}{\sigma_{p}},\quad$ (6)
$\text{MDD}=\max_{t=1}^{N}\left(\frac{P^{t}-P_{trough}^{t}}{P^{t}}\right)$ (7)
5.1.4. Implementation Details
@@
All LLM trading agents are deployed using GPT-4o Hurst et al. (2024) with a temperature setting of 0.7. All models are evaluated under the Position-Aware Trading Task formally defined in Sec. 3.2, ensuring profits and risks are calculated with explicit holding dynamics. Training covered Jan 2024–Feb 2025; testing spanned Mar–Apr 2025, coinciding with the U.S. election and other major macro events, which makes it a representative and challenging evaluation period.
Table 1. Performance comparison of different models on five stocks
@@
  MACD Wang and Kim (2018) 49.21-49.21 0.63-0.63 $74.20$ 46.83-46.83 0.57-0.57 $61.40$ 44.92-44.92 0.49-0.49 $70.30$ 7.42-7.42 0.18-0.18 $27.90$ 9.87-9.87 0.21-0.21 $19.80$
RSI Belafsky et al. (2002) 46.13-46.13 0.59-0.59 $72.80$ 44.28-44.28 0.54-0.54 $59.70$ 42.11-42.11 0.46-0.46 $68.20$ 6.98-6.98 0.16-0.16 $26.70$ 8.65-8.65 0.19-0.19 $19.30$
@@
  FinGPT Yang et al. (2023) 95.19-95.19 1.07-1.07 $94.36$ 85.22-85.22 1.52-1.52 $74.60$ 56.87-56.87 0.54-0.54 $70.30$ $2.01$ $0.01$ $22.30$ 3.63-3.63 0.04-0.04 $17.00$
FinAgent Zhang et al. (2024b) 74.31-74.31 0.89-0.89 $85.65$ 78.00-78.00 1.09-1.09 $71.50$ 49.64-49.64 0.53-0.53 $61.20$ $1.12$ $0.02$ $21.50$ 2.10-2.10 0.03-0.03 $17.50$
FINMEM Yu et al. (2024a) 44.03-44.03 0.52-0.52 $72.10$ 45.88-45.88 0.54-0.54 $47.80$ 41.21-41.21 0.36-0.36 $66.10$ $0.41$ $0.05$ $25.20$ $0.13$ $0.01$ $19.20$
FINCON Yu et al. (2024b) $7.76$ $0.38$ $59.13$ 16.02-16.02 0.13-0.13 $29.03$ 3.30-3.30 0.02-0.02 $54.77$ $7.92$ $0.63$ $20.70$ $7.35$ 1.03 $15.10$
@@
5.2. Main Results
We evaluated FinPos on five representative stocks—Tesla Inc. (TSLA, electric vehicles), Apple Inc. (AAPL, consumer technology), Amazon Inc. (AMZN, e-commerce), Netflix Inc. (NFLX, online entertainment), and Coinbase Global, Inc. (COIN, cryptocurrency trading). These assets span diverse sectors and exhibit distinct volatility characteristics during the test period: TSLA experienced the highest turbulence, AAPL and AMZN showed relatively stable dynamics, while NFLX and COIN displayed only mild fluctuations.
As shown in Tab 1, FinPos consistently outperforms all baselines, achieving higher CR, superior SR, and lower MDD across all five stocks. For high-volatility assets such as TSLA, AAPL, and AMZN, FinPos demonstrates remarkable robustness. While traditional DRL baselines such as A3C, DQN, and PPO all generated negative returns (e.g., A3C lost over 90% on TSLA with an MDD of 96.6%), and other LLM agents like FinGPT and FinAgent also suffered severe drawdowns (both exceeding 85% on TSLA), FinPos consistently maintained positive performance. Specifically, FinPos achieved CRs of 54.99%, 60.28%, and 52.10% on TSLA, AAPL, and AMZN respectively, while keeping MDDs controlled at 42.34%, 18.44%, and 36.92%. These results show that FinPos not only avoided catastrophic losses faced by DRL and LLM baselines, but also delivered sustained growth exceeding 50% on average under turbulent market conditions. For relatively stable assets such as NFLX and COIN, FinPos also delivered clear advantages. On NFLX, while DRL baselines hovered around neutral or slightly negative returns (e.g., PPO: 1.72%, DQN: 0.78%, A3C: -5.24%), FinPos achieved a CR of 16.85% with an SR of 0.87 and an MDD limited to 20.05%. Similarly, on COIN, FinPos obtained a CR of 14.74% , more than doubling the performance of LLM baselines such as FinCON (7.35%), while keeping downside risk under control (MDD 15.10%). These results demonstrate that FinPos maintains efficiency and steady growth even in calmer markets, combining superior profitability with stable risk management. This directly demonstrates the effectiveness of FinPos’s position-aware and risk-sensitive design.
5.3. Ablation Studies
We evaluated the contribution of each FinPos component through ablation experiments on three representative assets (TSLA, AAPL, AMZN).
  • Position Awareness (PA): Injects explicit position awareness into the agent by equipping the trading agent with a Quantity Decision module. Removing it turns the model into a position-agnostic agent, unable to condition its actions on current holdings.
  • Market Signal Processing (MSP): This module consists of multiple agents, each responsible for structuring and filtering raw data from different sources—including financial reports, news, and macroeconomic indicators—to provide coherent and refined market insights.
  • Multi-Timescale Reward (MTR): A self-reflection mechanism that enforces decision consistency across timescales without relying on external reward models.
We select TSLA, AAPL, and AMZN as test assets due to their rich set of news and events. Tab. 2 reports results under different module removals, showing that the full model consistently achieves the best CR, SR, and MDD. This demonstrates that FinPos’s various modules operate synergistically to yield optimal trading decisions.
Table 2. Ablation Results on TSLA, AAPL and AMZN
PA FIP MSP MTR TSLA AAPL AMZN
@@ 5.3. Ablation Studies
39.19 0.45 82.89 45.96 0.49 36.10 36.98 0.43 73.71
41.57 0.49 48.78 48.57 0.54 21.24 39.38 0.50 41.76
44.35 0.55 53.27 50.64 0.59 24.25 42.19 0.56 47.57
48.68 0.59 59.07 57.36 0.66 25.73 47.24 0.60 51.50
54.99 0.67 42.34 60.28 0.69 18.44 52.10 0.64 36.92
5.3.1. Position Awareness (PA)
As introduced in Sec. 4.2, the Trading Decision Module equips FinPos with position awareness. The ablation results in Tab. 2 confirm its necessity: on TSLA, CR drops from 54.99% to 39.19%, SR falls from 0.67 to 0.45, and MDD nearly doubles from 42.34% to 82.89% once PA is removed. Similar degradations appear on AAPL and AMZN. These findings validate that PA is indispensable for translating direction forecasts into position-based execution, ensuring both profitability and stability.
Figure 3. We conducted a fine-grained analysis of agent prompts along the dimensions of eight characteristics and three levels of emphasis.
@@ 5.3.1. Position Awareness (PA)
Figure 4. The impact of varying the maximum timescale of the multi-timescale reward on various performance metrics.
5.3.2. Financial Insight Prompting (FIP)
As shown in Tab. 2, introducing FIP leads to a clear overall performance gain, consistently enhancing cumulative return across all assets (e.g., TSLA: 41.57% → 54.99%; AAPL: 48.57% → 60.28%). In our ablation hierarchy, FIP ranks second only to the position-aware mechanism in terms of its contribution to overall performance.
To investigate the impact of prompt design on the depth of agent finance insight, we divide the content of prompts into eight key dimensions: Role Setting, Task Breakdown, Depth Analysis, Multi-Dimensional Analysis, Logical Reasoning, Contextual Consideration, Actionability, and Clarity & Conciseness(the detailed explanations of these eight dimensions are provided in Appendix C.2.1). For each dimension, we design prompts with different levels of refinement—Low, Medium, and High—and conduct graded experiments. The results are shown in Fig. 4.
5.3.3. Market Signal Processing (MSP) Module
We compare FinPos with a variant that directly feeds unfiltered raw data into the analysis agents. Removing MSP results in a consistent decline across all performance metrics (see Tab. 2). For instance, on TSLA, the cumulative return drops from 54.99% to 44.35%, the Sharpe Ratio drops from 0.67 to 0.55, and the maximum drawdown increases from 42.34% to 53.27%. Similar degradations are observed on AAPL (CR: 60.28% → 50.64%; SR: 0.69 → 0.59; MDD: 18.44% → 24.25%) and AMZN (CR: 52.10% → 42.19%; SR: 0.64 → 0.56; MDD: 36.92% → 47.57%). These results indicate that without structured filtering, the downstream analysis agents struggle to identify key signals from noisy textual inputs, leading to delayed or inaccurate trading responses. By contrast, the full FinPos system—with MSP-enabled preprocessing—produces cleaner and more coherent insights, thereby enhancing both trading precision and overall robustness.
5.3.4. Multi-Timescale Reward (MTR)
Finpos self-reflects its decision based on returns across multiple time scales (Sec. 4.3). Our experiments show that MTR is crucial for the agent’s long-term risk control. Without it, the agent loses sensitivity to long-horizon risks, leading to a sharp deterioration in drawdown management. As shown in Tab. 2, on TSLA, CR drops from 54.99% to 48.68%, SR decreases from 0.67 to 0.59, and MDD surges from 42.34% to 59.07%—a clear sign that the lack of multi-timescale self-reflection severely weakens the agent’s ability to capture stable trends and risk control. Similar degradations are observed on AAPL (CR: 60.28% → 57.36%; MDD: 18.44% → 25.73%) and AMZN (CR: 52.10% → 47.24%; MDD: 36.92% → 51.50%).
The effectiveness of MTR depends on the timescale used to compute the overall reward. In our design, the MTR module incorporates three reward horizons (1, 7, and 30 days) to jointly capture immediate fluctuations, medium-term trends, and long-term trends. Among these, the long-term trends serves as the principal timescale that governs the agent’s reflection and strategic consistency. Therefore, our sensitivity study focuses on the long-term trends, varying its length to examine the trade-off between adaptiveness and stability. As shown in Fig. 4, very short horizons (7–14 days) yield weak performance, as the agent focuses excessively on local fluctuations without forming stable strategies. Extending the horizon improves consistency, with 30 days emerging as the optimal setting: CR and SR peak, while drawdowns are minimized. However, when the horizon exceeds 30 days, performance declines again—longer windows dilute the reward signal and hinder the reflection process, making the agent less adaptive to new market conditions. This sensitivity study highlights that MTR is most effective when balancing short-term responsiveness with long-term stability, and in our experiments, 30 days provides the best trade-off.
5.4. Risk Analysis
Beyond numerical results, Fig. 5 provides complementary evidence from profitability and risk control. The top-left panel reports the Calmar ratio (see Appendix B.2), which directly reflects risk-adjusted performance under extreme losses. A higher value thus indicates that the strategy achieves strong profitability while effectively controlling downside risk. Here, FinPos achieves the highest score, suggesting that its profitability is supported by effective drawdown management rather than excessive risk-taking. The bottom-left panel plots cumulative return (CR %) against risk-control ability (1 - MDD). FinPos is positioned at the upper-right frontier, combining high profitability with robust risk management, whereas most baselines cluster in regions of either low return or weak control. On the right, the time-series curves and the exposure-risk overlay further illustrate the critical role of position-aware management: around high-volatility events (vertical dashed lines), methods without PA suffer sharp exposure spikes and subsequent drawdowns, while FinPos is capable of detecting the onset of risks and implementing defensive maneuvers. Taken together, these three views consistently demonstrate that position awareness not only mitigates short-term shocks but also improves risk-adjusted outcomes over the entire horizon. Additional analyses for other stocks are provided in Appendix C.1
Figure 5. Profitability versus risk dynamics on TSLA. Top-left: Calmar ratio, capturing return relative to maximum drawdown (higher is better). Bottom-left: joint view of profitability (CR%) and risk-control (1 - MDD), indicating return–risk balance. Right: time-series comparison of cumulative return (bottom) and exposure risk (top) across major events. Vertical dashed lines mark the occurrence of major events.
6. Conclusion

### v2

@@ 4.3 Multi-Timescale Reward Design
$d_{t}$ and $q_{t}$ denote the direction and quantity of the purchasing decision at the current time step, respectively. Crucially, this reward is used only during training-time simulation, no future-dependent signals are accessed at test time.
It serves two complementary roles in guiding the dual-agent architecture: (1) Guiding the Direction Agent: The term $\texttt{pos}_{t}\times M_{t}$ provides positive reinforcement when the agent’s directional decision (buy/sell/hold) aligns with the multi-horizon trend $M_{t}$. This encourages the Direction Agent to move beyond short-term noise and base decisions on longer-term market signals. (2) Guiding the Quantity Agent: During LLM reflection, the Quantity Agent uses this reward to learn when to adjust exposure. Through this process, the agent internalizes the concept of position as a dynamic state. In our experiments, we find that during periods of high volatility, agents tend to maintain a inactive state. To address this, we introduce a quadratic penalty on $|M_{t}|$. whenever the position remains unchanged. This discourages missed opportunities in volatile markets while also preventing excessive trading in stable periods.
5 Experiments
5.1 Experimental Setup
5.1.1 Datasets
@@
Our research utilizes actual financial data sourced from publicly authoritative providers and encompasses various forms of market information. (1) Stock prices, obtained from Yahoo Finance, include daily open-high-low-close-volume (OHLCV) data. (2) Company news, retrieved using the Finnhub stock API, includes articles’ related company name, publication date, headline, and summary text, processed for sentiment and semantic relevance. (3) Macroeconomic news, also retrieved via the Finnhub stock API, contains publication dates, headlines, and summary texts for each article. (4) 10-Q (quarterly reports) and 10-K (annual reports), accessed through the U.S. Securities and Exchange Commission (SEC) EDGAR API. These documents provide financial information that publicly listed companies are required to disclose, and standardized into a daily time series format.
5.1.2 Comparative Methods
@@
To assess the effectiveness of FinPos, we conduct comparative experiments with four LLM-based trading agents: FinGPT Yang et al. (2023), FINMEM Yu et al. (2024a), FinAgent Zhang et al. (2024), and FINCON Yu et al. (2024b); three deep reinforcement learning (DRL) methods (A2C, PPO, DQN); two rule-based methods (MACD, RSI); and two market baselines (Random). All compared models are evaluated using identical data splits and executed under the same market conditions. Additional details are provided in Appendix C.
5.1.3 Evaluation Metrics
@@
We evaluate performance using cumulative return (CR%), sharpe ratio (SR), and maximum drawdown (MDD%). CR reflects overall profitability, while SR measures the risk-adjusted return by showing how much excess return it generates per unit of total risk. MDD quantifies the worst-case loss over the trading horizon, which in the position-aware trading task directly reflects exposure and vulnerability under held positions. CR has been introduced earlier (see Eq. (2)), and the formulas for SR and MDD are provided in Appendix B.1.
5.1.4 Implementation Details
@@
All LLM trading agents are deployed using GPT-4o Hurst et al. (2024) with a temperature setting of 0.7. All models are evaluated under the position-aware trading task formally defined in Sec. 3.2, ensuring profits and risks are calculated with explicit holding dynamics. Training covered Jan 2024–Feb 2025; testing spanned Mar–Sep 2025, covering the U.S. election period and other major macroeconomic events, making it a representative and challenging evaluation setting.
Table 1: Performance comparison of different models on five stocks
@@
  MACD Wang and Kim (2018) 46.25-46.25 0.55-0.55 $74.20$ 29.61-29.61 0.42-0.42 $51.40$ 38.14-38.14 0.47-0.47 $70.30$ 27.45-27.45 0.23-0.23 $47.92$ 49.87-49.87 0.54-0.54 $70.83$
RSI Belafsky et al. (2002) 45.06-45.06 0.52-0.52 $72.80$ 33.28-33.28 0.48-0.48 $49.70$ 35.10-35.10 0.42-0.42 $68.20$ 26.98-26.98 0.21-0.21 $46.08$ 53.65-53.65 0.63-0.63 $66.37$
@@
  FinGPT Yang et al. (2023) 89.36-89.36 0.98-0.98 $94.36$ 80.37-80.37 1.01-1.01 $74.60$ 60.14-60.14 0.67-0.67 $70.30$ 26.01-26.01 0.40-0.40 $57.30$ 42.63-42.63 0.34-0.34 $67.09$
FinAgent Zhang et al. (2024) 65.07-65.07 0.76-0.76 $85.65$ 51.10-51.10 0.69-0.69 $49.50$ 43.29-43.29 0.47-0.47 $61.20$ 18.71-18.71 0.22-0.22 $41.50$ $2.76$ $0.23$ $56.50$
FINMEM Yu et al. (2024a) 36.48-36.48 0.45-0.45 $72.10$ 38.94-38.94 0.52-0.52 $44.82$ 27.61-27.61 0.33-0.33 $66.10$ 33.75-33.75 0.37-0.37 $60.02$ 2.13-2.13 0.12-0.12 $51.28$
FINCON Yu et al. (2024b) $19.67$ $0.36$ $59.13$ 2.81-2.81 0.06-0.06 $31.03$ $7.96$ $0.12$ $34.77$ $6.12$ $0.17$ $34.70$ $21.35$ $0.63$ $46.10$
@@
5.2 Main Results
We evaluated FinPos on five representative stocks: TSLA, AAPL, AMZN, NFLX, and COIN. As shown in Tab. 1, FinPos consistently outperforms all baseline methods, achieving higher CRs, SRs, and lower MDDs across all assets. On high-volatility stocks such as TSLA, AAPL, and COIN, FinPos exhibits strong robustness under turbulent market conditions. While DRL baselines (A3C, DQN, PPO) and LLM-based agents (FinGPT, FinAgent) suffer substantial losses and severe drawdowns (e.g., A3C loses over 80% on TSLA with an MDD of 96.6%), FinPos consistently maintains positive returns, achieving CRs of 62.15% and 36.31% on TSLA and AAPL, respectively, with well-controlled downside risk. For assets with clearer trends, such as AMZN and NFLX, FinPos continues to deliver stable and risk-adjusted performance. FinPos achieves a CR of 30.35% on AMZN with an MDD of 18.44%, and a CR of 28.65% on NFLX with a high SR of 1.02, outperforming all baseline agents in both return and risk control.
5.3 Ablation Studies
We select TSLA, AAPL, and AMZN as test assets due to their rich coverage of news articles, earnings events, and macroeconomic sensitivities. We assess the contribution of each FinPos component through ablation experiments. Tab. 2 reports results under different module removals, showing that the full model consistently achieves the best performance. Extended ablation results are provided in Appendix D.3.
  • Multi-Timescale Reward (MTR) is the cornerstone of our position-aware framework. It provides a self-supervised signal that aligns short-term actions with long-horizon outcomes, enabling the agent to internalize the cumulative impact of position decisions.
  • Quantity and Risk Decision Agent (QRA) is essential for explicit position management. It transforms the agent from a unit-action trader into a continuous-position controller. Removing QRA forces fixed-unit trades, eliminating dynamic exposure control.
  • Market Signal Processing (MSP) acts as a preprocessing filter that scores raw market data by relevance and importance before downstream agents analyze them, ensuring that only high-signal information enters the decision pipeline.
Table 2: Ablation Results on TSLA, AAPL and AMZN
MTR QRA MSP TSLA AAPL AMZN
@@ 5.3 Ablation Studies
18.73 0.35 57.87 12.93 0.20 35.63 9.75 0.21 26.20
53.57 0.49 62.65 29.30 0.31 39.29 27.85 0.43 30.29
58.34 0.63 45.40 34.09 0.39 29.87 28.50 0.32 19.87
62.15 0.68 42.34 36.31 0.43 27.53 30.35 0.34 18.44
@@
Figure 3: Impact of varying the maximum timescale of the multi-timescale reward on performance metrics.
5.3.1 Multi-Timescale Reward (MTR)
As shown in Tab. 2, removing MTR causes a severe performance drop across all assets: the CR fall below 20% for all three stocks (vs. 30.4–62.2% in the full model), indicating that MTR is the architectural backbone that enables long-horizon reasoning and strategic consistency in FinPos.
The effectiveness of MTR depends critically on the choice of timescale for computing the cumulative reward. Our design incorporates three horizons (1, 7, and 30 days) to capture immediate fluctuations, medium-term trends, and long-term stability, respectively. We conduct a sensitivity study by varying the length of the long-term horizon (Fig. 3), finding that: very short windows (7–14 days) yield poor performance, as the agent overreacts to noise without forming stable strategies; performance peaks at 30 days; beyond 30 days, performance declines due to signal dilution, which weakens the reflection mechanism and reduces adaptiveness to regime shifts.
5.3.2 Quantity and Risk Decision Agent (QRA)
Removing QRA leads to a significant degradation in risk control. As shown in Tab. 2, while CR drop moderately (e.g., from 62.2% to 53.6% on TSLA), the MDD worsens dramatically—increasing from 42.3% to 62.7% on TSLA (a 48% relative increase) and from 27.5% to 39.3% on AAPL. This reveals a critical limitation: even with long-horizon awareness provided by MTR, an agent restricted to fixed-unit trades cannot timely adjust its exposure in response to market volatility. Consequently, it lacks the agility needed to hedge positions during sharp, fast-moving drawdowns—precisely the scenarios where position-aware sizing is most valuable.
5.3.3 Market Signal Processing (MSP)
Removing MSP leads to moderate performance degradation (e.g., on TSLA: CR drops from 62.2% to 58.3%, and MDD worsens from 42.3% to 45.4%; see Tab. 2), revealing that raw market data contain substantial noise that distracts downstream agents. Without MSP’s relevance filtering, an agent may erroneously treat Elon Musk’s tweets about Dogecoin as material signals for TSLA trading, despite their lack of fundamental relevance. Such off-topic social media noise is common in financial news feeds and can trigger spurious trading responses. In contrast, MSP leverages domain-guided heuristics and entity-relevance rules to pre-filter inputs, ensuring that only high-signal information—such as regulatory filings, or company-specific announcements—reaches the LLM analysis pipeline.
5.4 Risk Analysis
Beyond numerical results, Fig. 4 provides focused evidence of FinPos’s risk-adjusted performance during a high-volatility period (Mar–Apr 2025), chosen to highlight its ability to manage extreme market conditions—including those triggered by major political events such as the U.S. election cycle. This window offers a stringent test for position-awareness, as agents must adapt rapidly to shifting risk regimes without overreacting. The top-left panel reports the calmar ratio (Appendix B.3), which measures return per unit of maximum drawdown. FinPos achieves a calmar of 1.5, significantly outperforming all baselines, indicating that its profitability is not driven by excessive risk-taking but by effective control. The bottom-left panel plots cumulative return (CR%) against risk-control ability (1-MDD). FinPos occupies the upper-right frontier—high return with low drawdown—while baselines cluster in regions of either low return or poor risk management. On the right, the time-series curves and exposure-risk overlay reveal the operational advantage of position-aware management: during high-volatility events (marked by vertical dashed lines), non-PA agents exhibit sharp exposure spikes followed by severe drawdowns, while FinPos proactively reduces exposure and stabilizes returns. This demonstrates that position awareness enables not only reactive mitigation but also anticipatory adjustment—critical for navigating real-world market turbulence. Additional analyses for AAPL are provided in Appendix D.2.
Figure 4: Risk-adjusted performance and exposure dynamics on TSLA during Mar–Apr 2025, highlighting FinPos’s advantage under high-volatility events.
6 Conclusion